# Counting rational maps on $\mathbb{P}^{1}$ with prescribed local conditions

### v2

@@ 3. Projective sets with prescribed local conditions and their affine liftings
$\{(x_{0},\dots,x_{m})\in R^{m+1}:\ \text{$x_{i}$ is a unit for some $i$}\}$
@@ 3.1. Local conditions and affine liftings
(3.1) $\pi(\mathcal{S}_{p}):=\frac{\#\bar{\mathcal{S}}_{p}}{\#\mathbb{P}^{2d+1}\left(\mathbb{Z}/p^{n_{p}}\mathbb{Z}\right)}=\frac{p^{n_{p}-1}(p-1)\#\bar{\mathcal{S}}_{p}}{p^{(2d+2)n_{p}}\left(1-p^{-(2d+2)}\right)}.$
@@
From #$\#\bar{S}_{p}=\#(\mathbb{Z}/p^{n_{p}}\mathbb{Z})^{*}\cdot\#\bar{\mathcal{S}}_{p}=p^{n_{p}-1}(p-1)\#\bar{\mathcal{S}}_{p}$, we have
@@
Finally, if $\mathcal{S}=\bigcap_{p}\mathcal{S}_{p}$ is defined by local conditions then the affine lifting $S$ of $\mathcal{S}$ is the set $S=\bigcap_{p}S_{p}$ where each $S_{p}$ is the affine lifting of $\mathcal{S}_{p}$. It is immediate from the definition that $S$ consists precisely the points $Q=(C_{1},\ldots,C_{2d+2})$ such that $\gcd(C_{1},\ldots,C_{2d+2})=1$ and $[Q]\in\mathcal{S}_{p}$ for every $p$. In other words, from Notation 2.2, $S$ consists of exactly pairs of points $Q$ and $-Q$ such that one of them belongs to $\mathcal{S}$. Hence for every $\vec{r}$, we have:
(3.5) #$\#\mathcal{S}(\vec{r})=\frac{1}{2}\#S(\vec{r}).$
@@
$\lim_{Z\rightarrow\infty}\limsup_{r_{1},\ldots,r_{2d+2}\rightarrow\infty}\frac{\#T_{Z}(\vec{r})}{\#\operatorname{Box}(\vec{r})}=0.$
@@
#$\#T_{Z}(\vec{r})=O\left(\sum_{p>Z}\frac{r_{1}\cdots r_{2d+2}}{p^{2d+2}}\right)=O\left(\frac{r_{1}\cdots r_{2d+2}}{Z^{2d+1}}\right)$
@@
$S_{p}=\{(C_{1},\ldots,C_{2d+2})\in\mathbb{Z}^{2d+2}:\ p\nmid C_{i}\ \text{for some $i$}\}.$
Now $\mathcal{S}:=\bigcap_{p}\mathcal{S}_{p}=\mathbb{P}^{2d+1}(\mathbb{Q})$ and its affine lifting is $S:=\bigcap_{p}S_{p}$.
@@
Recall the set $T_{>Z}$ in Lemma 3.1. From the definitions of the involving sets, we have:
@@ 3.2. Some admissibility conditions
We interpret $\prod_{p}\pi(\mathcal{S}_{p})$ as the limit of the sequence of partial products $\prod_{p\leq k}\pi(\mathcal{S}_{p})$. This limit exists (with value in $[0,1]$) since $0\leq\pi(\mathcal{S}_{p})\leq 1$ for every $p$. The same holds for $\prod_{p}\pi(S_{p})$. Our admissibility condition is analogous to the admissibility condition for Weierstrass models considered by Cremona and Sadek in [CS23, Definition 4].
@@
    The first consists of $Q=(C_{1},\ldots,C_{2d+2})$ such that $\gcd(C_{1},\ldots,C_{2d+2})=1$. These points come in pairs (namely $Q$ and $-Q$) and, according to Notation 2.2 and the definitions of the involving sets, exactly one point in each pair belongs to $\mathcal{S}_{>Z}^{\prime}$.
@@
#$\#(S_{\leq Z}\cap S_{>Z}^{\prime})(\vec{r})\leq 2\#\mathcal{S}_{>Z}^{\prime}(\vec{r})+\#T_{>Z}(\vec{r}).$
@@ 4. Minimal resultants
In this section, fix an integer $d\geq 2$ and our goal is to prove that a positive portion of $\mathcal{R}^{(d)}$ has globally minimal resultants. In fact, we consider a certain subset of $\mathcal{R}^{(d)}$ that includes globally minimal rational maps that have good reduction at a given finite set of primes and proves that this subset has positive height density (for odd $d$) as well as positive weak box density (for any $d$). In section 5, we obtain an explicit lower bound for the weak box density in the case $d=2$.
@@ 4.1. The minimal resultant
Let $p$ be a prime number and set $\widetilde{F}(X,Y)$ and $\widetilde{G}(X,Y)$ to denote the mod-$p$ reductions of $F(X,Y)$ and $G(X,Y)$ respectively, and set $\tilde{\phi}(X/Y):=\frac{\widetilde{F}(X,Y)}{\widetilde{G}(X,Y)}$. As is well known [Sil07, Theorem 2.15], the following conditions are equivalent
@@
Let $\mathrm{ord}_{p}$ denote the usual $p$-adic valuation on $\mathbb{Q}$. The following definition is taken from [Sil07, Chapter 4].
@@
  • Given $A\in\operatorname{PGL}_{2}(\mathbb{Q})$, set $\phi^{A}:=A^{-1}\circ\phi\circ A$. We set $\epsilon_{p}(\phi)$ to be the minimum value of $\operatorname{ord}_{p}(\operatorname{Res}(\phi^{A}))$, where $A$ ranges over $\operatorname{PGL}_{2}(\mathbb{Q})$.
@@
    $\mathfrak{R}_{\phi}:=\prod_{p}p^{\epsilon_{p}(\phi)}.$
@@
The result is a special case of [BM12, Lemma 3.1]. ∎
We now introduce the relevant subsets of $\mathcal{R}^{(d)}$ defined by local conditions. Let $\mathcal{N}_{p}^{(k)}$ consist of $\phi\in\mathcal{R}^{(d)}$ such that $p^{k}\nmid\operatorname{Res}(\phi)$ so that $\mathcal{G}_{p}:=\mathcal{N}_{p}^{(1)}$ is the set of rational maps with good reduction at $p$.
@@
$\begin{split}&\mathcal{V}_{\Sigma}:=\left(\bigcap_{p\in\Sigma}\mathcal{G}_{p}\right)\cap\left(\bigcap_{p\notin\Sigma}\mathcal{N}_{p}^{(d)}\right);\\ &\mathcal{W}_{\Sigma}:=\left(\bigcap_{p\in\Sigma}\mathcal{G}_{p}\right)\cap\left(\bigcap_{p\notin\Sigma}\mathcal{N}_{p}^{(2d)}\right).\end{split}$
@@ 4.2. Equations over finite fields
$\bar{\mathcal{G}}_{p}:=\{P\in\mathbb{P}^{2d+1}(\mathbb{Z}/p\mathbb{Z}):\ \mathfrak{r}(P)\neq 0\}$
@@
$\bar{G}_{p}:=\{Q\in(\mathbb{Z}/p\mathbb{Z})^{2d+2}:\ \mathfrak{r}(Q)\neq 0\}.$
The preimage $G_{p}$ of $\bar{G}_{p}$ under the reduction map $\mathbb{Z}^{2d+2}\rightarrow(\mathbb{Z}/p\mathbb{Z})^{2d+2}$ is the affine lifting of $\mathcal{G}_{p}$. The complement $\bar{G}_{p}^{\prime}:=(\mathbb{Z}/p\mathbb{Z})^{2d+2}\setminus\bar{G}_{p}$ is the zero set of $\mathfrak{r}$ in $(\mathbb{Z}/p\mathbb{Z})^{2d+2}$. By (3.4), we have
@@
We note that #$\#\bar{G}_{p}^{\prime}=p^{2d+2}-\#\bar{G}_{p}$. Then part (1) follows from Theorem 4.5. For part (2), the lower bound #$\#\bar{G}_{p}^{\prime}\geq p^{2d+2-2d}=p^{2}$ follows from Theorem 4.6. Next, part (3) follows from Theorem 4.7. Finally for part (4), since $Y^{d}$ and $X^{d}$ have no common root in $\mathbb{F}_{p}$, we have that $\mathfrak{r}(1,0,\ldots,0,1)=:c$ is a non-zero element of $\mathbb{F}_{p}$. Then we apply Theorem 4.6 for the polynomial $\mathfrak{r}-c$. ∎
@@ 4.3. Some positive densities
$\bar{\mathcal{N}}_{p}^{(k)}:=\{P\in\mathbb{P}^{2d+1}(\mathbb{Z}/p^{k}\mathbb{Z}):\ \mathfrak{r}(P)\neq 0\ \text{in}\ \mathbb{Z}/p^{k}\mathbb{Z}\}$
@@
$\bar{N}_{p}^{(k)}:=\{Q\in(\mathbb{Z}/p^{k}\mathbb{Z})^{2d+2}:\ \mathfrak{r}(Q)\neq 0\ \text{in}\ \mathbb{Z}/p^{k}\mathbb{Z}\}.$
@@
The complement $(\mathcal{N}_{p}^{(k)})^{\prime}:=\mathbb{P}^{2d+1}(\mathbb{Q})\setminus\mathcal{N}_{p}^{(k)}$ is defined by a local condition since it is the preimage of
$(\bar{\mathcal{N}}_{p}^{(k)})^{\prime}:=\mathbb{P}^{2d+1}(\mathbb{Z}/p^{k}\mathbb{Z})\setminus\bar{\mathcal{N}}_{p}^{(k)}=\{P\in\mathbb{P}^{2d+1}(\mathbb{Z}/p^{k}\mathbb{Z}):\ \mathfrak{r}(P)=0\ \text{in}\ \mathbb{Z}/p^{k}\mathbb{Z}\}.$
@@
The complement $(\bar{N}_{p}^{(k)})^{\prime}:=(\mathbb{Z}/p^{k}\mathbb{Z})^{2d+2}\setminus\bar{N}_{p}^{(k)}$ is the zero set of $\mathfrak{r}$ in $(\mathbb{Z}/p^{k}\mathbb{Z})^{2d+2}$ and the complement $(N_{p}^{(k)})^{\prime}:=\mathbb{Z}^{2d+2}\setminus N_{p}^{(k)}$ is the preimage of $(\bar{N}_{p}^{(k)})^{\prime}$ under the reduction map $\mathbb{Z}^{2d+2}\rightarrow(\mathbb{Z}/p^{k}\mathbb{Z})^{2d+2}$. As in Subsection 3.2, we warn the reader that neither $(\bar{N}_{p}^{(k)})^{\prime}$ is the affine lifting of $(\bar{\mathcal{N}}_{p}^{(k)})^{\prime}$ nor $(N_{p}^{(k)})^{\prime}$ is the affine lifting of $(\mathcal{N}_{p}^{(k)})^{\prime}$ (roughly speaking, this means taking the complement and taking the affine lifting do not commute to each other). In fact, the affine lifting of $(\bar{\mathcal{N}}_{p}^{(k)})^{\prime}$ is the set of points in $(\bar{N}_{p}^{(k)})^{\prime}$ having at least one entry that is a unit of $\mathbb{Z}/p^{k}\mathbb{Z}$. Similarly, the affine lifting of $(\mathcal{N}_{p}^{(k)})^{\prime}$ is the set of points in $(N_{p}^{(k)})^{\prime}$ having at least one entry that is not divisible by $p$. Therefore:
@@
In order to obtain an upper bound for #$\#(\bar{N}_{p}^{(k)})^{\prime}$, we use the following result
@@
This is proved in [GKZ08]. ∎
@@
For the rest of this section, let $\mathfrak{s}$ denote one of the partial derivatives of $\mathfrak{r}$ (when necessary, which partial derivative will be specified precisely in later context). This is a homogeneous polynomial of degree $12d-$ 1 and has degree $d-1$ in the variable for which we differentiate.
@@
#$\#\{P\in\mathbb{F}_{p}^{2d+2}\mid\mathfrak{r}(P)=\mathfrak{s}(P)=0\}\leq\gamma_{1}p^{2d}$.
@@
Consider the algebra $\mathfrak{B}:=\mathbb{Q}[x_{0},\dots,x_{d},y_{0},\dots,y_{d}]/(\mathfrak{r},\mathfrak{s})$ and let $\bar{x}_{i}$ and $\bar{y}_{j}$ denote the image of $x_{i}$ and $y_{j}$ in $\mathfrak{B}$. Since $\mathfrak{r}$ is irreducible and does not divide $\mathfrak{s}$, the dimension of $\mathfrak{B}$ is $D\leq 2d$. By the Noether normalization lemma, there exist $X_{1},\ldots,X_{D}\in\mathfrak{B}$ that are algebraically independent over $\mathbb{Q}$ such that $\mathfrak{B}$ is integral over $\mathbb{Q}[X_{1},\dots,X_{D}]$. Thus there is a natural number $N$ such that the $\bar{x}_{i}$’s and $\bar{y}_{j}$’s are all integral over $\mathbb{Z}[1/N,X_{1},\dots,X_{D}]$. Suppose $p\nmid N$, we find that any tuple $(X_{1},\dots,X_{D})\in(\mathbb{Z}/p\mathbb{Z})^{D}$ gives rise to at most $O(1)$ many
$P:=(A_{0},\dots,A_{d},B_{0},\dots,B_{d})\in(\mathbb{Z}/p\mathbb{Z})^{2d+2}$
with $\mathfrak{r}(P)=\mathfrak{s}(P)=0$. If we also require $\gamma_{1}\geq N^{2}$ then the inequality in the statement of the proposition holds trivially for $p\mid N$. This finishes the proof. ∎
@@
Each element of $S_{1}$ lifts to $p^{2d+2}$ many elements of $(\mathbb{Z}/p^{2}\mathbb{Z})^{2d+2}$ and we have #$\#S_{1}\leq\gamma_{1}p^{2d}$ by Proposition 4.13. Each element of $S_{2}$ lifts to $p^{2d+1}$ many elements $\beta\in(\mathbb{Z}/p^{2}\mathbb{Z})^{2d+2}$ such that $\mathfrak{r}(\beta)=0\bmod p^{2}$ by Lemma 4.15 and we have #$\#S_{2}\leq 2dp^{2d+1}$ by Theorem 4.7. Overall, we have:
#$\#(\mathcal{N}_{p}^{(2)})^{\prime}\leq(\gamma_{1}+2d)p^{4d+2}.$
@@
The set $W_{\Sigma}:=\left(\bigcap_{p\in\Sigma}G_{p}\right)\cap\left(\bigcap_{p\notin\Sigma}N_{p}^{(2d)}\right)$ is the affine lifting of $\mathcal{W}_{\Sigma}$. Recall the complement $(N_{p}^{(2d)})^{\prime}=\mathbb{Z}^{2d+2}\setminus N_{p}^{(2d)}$. We recall the notation at the beginning of Subsection 3.2. Let $Z$ be a positive real number such that $Z>\max\{p\in\Sigma\}$ and let
@@
$T_{>Z}(x):=W_{>Z}^{\prime}(x)\setminus S_{>Z}(x).$
It follows from [Poo03, Lemma 5.1] that
@@
$h_{Q}(u):=\mathfrak{r}(u,A_{1},\dots,A_{d},B_{0},\dots,B_{d-1},B_{d})\in\mathbb{Z}[u],$

### v3

@@ 3. Projective sets with prescribed local conditions and their affine liftings
$\{(x_{0},\dots,x_{m})\in R^{m+1}\mathrel{\mathop{\ordinarycolon}}\ \text{$x_{i}$ is a unit for some $i$}\}$
@@ 3.1. Local conditions and affine liftings
(3.1) $\pi(\mathcal{S}_{p})\mathrel{\mathop{\ordinarycolon}}=\frac{\#\bar{\mathcal{S}}_{p}}{\#\mathbb{P}^{2d+1}\left(\mathbb{Z}/p^{n_{p}}\mathbb{Z}\right)}=\frac{p^{n_{p}-1}(p-1)\#\bar{\mathcal{S}}_{p}}{p^{(2d+2)n_{p}}\left(1-p^{-(2d+2)}\right)}.$
@@
From $\bar{S}_{p}=\#(\mathbb{Z}/p^{n_{p}}\mathbb{Z})^{*}\cdot\#\bar{\mathcal{S}}_{p}=p^{n_{p}-1}(p-1)\#\bar{\mathcal{S}}_{p}$, we have
@@
Finally, if $\mathcal{S}=\bigcap_{p}\mathcal{S}_{p}$ is defined by local conditions then the affine lifting $S$ of $\mathcal{S}$ is the set $S=\bigcap_{p}S_{p}$ where each $S_{p}$ is the affine lifting of $\mathcal{S}_{p}$. It is immediate from the definition that $S$ consists precisely the points $Q=(C_{1},\ldots,C_{2d+2})$ such that $\gcd(C_{1},\ldots,C_{2d+2})=1$ and $Q]\in\mathcal{S}_{p}$ for every $p$. In other words, from Notation 2.2, $S$ consists of exactly pairs of points $Q$ and Q-Q such that one of them belongs to $\mathcal{S}$. Hence for every $\vec{r}$, we have:
(3.5) $\mathcal{S}(\vec{r})=\frac{1}{2}\#S(\vec{r}).$
@@
$\lim_{Z\rightarrow\infty}\limsup_{r_{1},\ldots,r_{2d+2}\rightarrow\infty}\frac{\#T_{>Z}(\vec{r})}{\#\operatorname{Box}(\vec{r})}=0.$
@@
$T_{>Z}(\vec{r})=O\left(\sum_{p>Z}\frac{r_{1}\cdots r_{2d+2}}{p^{2d+2}}\right)=O\left(\frac{r_{1}\cdots r_{2d+2}}{Z^{2d+1}}\right)$
@@
$S_{p}=\{(C_{1},\ldots,C_{2d+2})\in\mathbb{Z}^{2d+2}\mathrel{\mathop{\ordinarycolon}}\ p\nmid C_{i}\ \text{for some $i$}\}.$
Now $\mathcal{S}\mathrel{\mathop{\ordinarycolon}}=\bigcap_{p}\mathcal{S}_{p}=\mathbb{P}^{2d+1}(\mathbb{Q})$ and its affine lifting is $S\mathrel{\mathop{\ordinarycolon}}=\bigcap_{p}S_{p}$.
@@
Recall the set $T_{>Z}$ in Lemma 3.1. From the definitions of the involved sets, we have:
@@ 3.2. Some admissibility conditions
We interpret $\prod_{p}\pi(\mathcal{S}_{p})$ as the limit of the sequence of partial products $\prod_{p\leq k}\pi(\mathcal{S}_{p})$. This limit exists (with value in $[0,1]$) since $0\leq\pi(\mathcal{S}_{p})\leq 1$ for every $p$. The same holds for $\prod_{p}\pi(S_{p})$. Our admissibility condition is analogous to the admissibility condition for Weierstrass models considered by Cremona and Sadek in [cremonasadek, Definition 4].
@@
    The first consists of $Q=(C_{1},\ldots,C_{2d+2})$ such that $\gcd(C_{1},\ldots,C_{2d+2})=1$. These points come in pairs (namely $Q$ and Q-Q) and, according to Notation 2.2 and the definitions of the involving sets, exactly one point in each pair belongs to $\mathcal{S}_{>Z}^{\prime}$.
@@
$(S_{\leq Z}\cap S_{>Z}^{\prime})(\vec{r})\leq 2\#\mathcal{S}_{>Z}^{\prime}(\vec{r})+\#T_{>Z}(\vec{r}).$
@@ 4. Minimal resultants
In this section, fix an integer $d\geq 2$ and our goal is to prove that a positive portion of $\mathcal{R}^{(d)}$ has globally minimal resultants. In fact, we consider a certain subset of $\mathcal{R}^{(d)}$ that includes globally minimal rational maps that have good reduction at a given finite set of primes and prove that this subset has positive height density (for odd $d$) as well as positive weak box density (for any $d$). In section 5, we obtain an explicit lower bound for the weak box density in the case $d=2$.
@@ 4.1. The minimal resultant
Let $p$ be a prime number and set $\widetilde{F}(X,Y)$ and $\widetilde{G}(X,Y)$ to denote the mod-$p$ reductions of $F(X,Y)$ and $G(X,Y)$ respectively, and set $\tilde{\phi}(X/Y)\mathrel{\mathop{\ordinarycolon}}=\frac{\widetilde{F}(X,Y)}{\widetilde{G}(X,Y)}$. As is well known [silvermantextbook, Theorem 2.15], the following conditions are equivalent
@@
Let $\mathrm{ord}_{p}$ denote the usual $p$-adic valuation on $\mathbb{Q}$. The following definition is taken from [silvermantextbook, Chapter 4].
@@
  • Given $A\in\operatorname{PGL}_{2}(\mathbb{Q})$, set $\phi^{A}\mathrel{\mathop{\ordinarycolon}}=A^{-1}\circ\phi\circ A$. We set $\epsilon_{p}(\phi)$ to be the minimum value of $\operatorname{ord}_{p}(\operatorname{Res}(\phi^{A}))$, where $A$ ranges over $\operatorname{PGL}_{2}(\mathbb{Q})$.
@@
    $\mathfrak{R}_{\phi}\mathrel{\mathop{\ordinarycolon}}=\prod_{p}p^{\epsilon_{p}(\phi)}.$
@@
The result follows immediately from [silvermantextbook, Proposition 4.95]. ∎
We now introduce the relevant subsets of $\mathcal{R}^{(d)}$ defined by local conditions. Let $\mathcal{N}_{p}^{(k)}$ consist of $\phi\in\mathcal{R}^{(d)}$ such that $p^{k}\nmid\operatorname{Res}(\phi)$ so that $\mathcal{G}_{p}\mathrel{\mathop{\ordinarycolon}}=\mathcal{N}_{p}^{(1)}$ is the set of rational maps with good reduction at $p$.
@@
$\begin{split}&\mathcal{V}_{\Sigma}\mathrel{\mathop{\ordinarycolon}}=\left(\bigcap_{p\in\Sigma}\mathcal{G}_{p}\right)\cap\left(\bigcap_{p\notin\Sigma}\mathcal{N}_{p}^{(d)}\right);\\ &\mathcal{W}_{\Sigma}\mathrel{\mathop{\ordinarycolon}}=\left(\bigcap_{p\in\Sigma}\mathcal{G}_{p}\right)\cap\left(\bigcap_{p\notin\Sigma}\mathcal{N}_{p}^{(2d)}\right).\end{split}$
@@ 4.2. Equations over finite fields
$\bar{\mathcal{G}}_{p}\mathrel{\mathop{\ordinarycolon}}=\{P\in\mathbb{P}^{2d+1}(\mathbb{Z}/p\mathbb{Z})\mathrel{\mathop{\ordinarycolon}}\ \mathfrak{r}(P)\neq 0\}$
@@
$\bar{G}_{p}\mathrel{\mathop{\ordinarycolon}}=\{Q\in(\mathbb{Z}/p\mathbb{Z})^{2d+2}\mathrel{\mathop{\ordinarycolon}}\ \mathfrak{r}(Q)\neq 0\}.$
The preimage $G_{p}$ of $\bar{G}_{p}$ under the reduction map $\mathbb{Z}^{2d+2}\rightarrow(\mathbb{Z}/p\mathbb{Z})^{2d+2}$ is the affine lifting of $\mathcal{G}_{p}$. The complement $\bar{G}_{p}^{\prime}\mathrel{\mathop{\ordinarycolon}}=(\mathbb{Z}/p\mathbb{Z})^{2d+2}\setminus\bar{G}_{p}$ is the zero set of $\mathfrak{r}$ in $(\mathbb{Z}/p\mathbb{Z})^{2d+2}$. By (3.4), we have
@@
We note that $\bar{G}_{p}^{\prime}=p^{2d+2}-\#\bar{G}_{p}$. Then part (1) follows from Theorem 4.5. For part (2), the lower bound $\bar{G}_{p}^{\prime}\geq p^{2d+2-2d}=p^{2}$ follows from Theorem 4.6. Next, part (3) follows from Theorem 4.7. Finally for part (4), since $Y^{d}$ and $X^{d}$ have no common root in $\mathbb{F}_{p}$, we have that $\mathfrak{r}(1,0,\ldots,0,1)=\mathrel{\mathop{\ordinarycolon}}c$ is a non-zero element of $\mathbb{F}_{p}$. Then we apply Theorem 4.6 for the polynomial $\mathfrak{r}-c$. ∎
@@ 4.3. Some positive densities
$\bar{\mathcal{N}}_{p}^{(k)}\mathrel{\mathop{\ordinarycolon}}=\{P\in\mathbb{P}^{2d+1}(\mathbb{Z}/p^{k}\mathbb{Z})\mathrel{\mathop{\ordinarycolon}}\ \mathfrak{r}(P)\neq 0\ \text{in}\ \mathbb{Z}/p^{k}\mathbb{Z}\}$
@@
$\bar{N}_{p}^{(k)}\mathrel{\mathop{\ordinarycolon}}=\{Q\in(\mathbb{Z}/p^{k}\mathbb{Z})^{2d+2}\mathrel{\mathop{\ordinarycolon}}\ \mathfrak{r}(Q)\neq 0\ \text{in}\ \mathbb{Z}/p^{k}\mathbb{Z}\}.$
@@
The complement $(\mathcal{N}_{p}^{(k)})^{\prime}\mathrel{\mathop{\ordinarycolon}}=\mathbb{P}^{2d+1}(\mathbb{Q})\setminus\mathcal{N}_{p}^{(k)}$ is defined by a local condition since it is the preimage of
$(\bar{\mathcal{N}}_{p}^{(k)})^{\prime}\mathrel{\mathop{\ordinarycolon}}=\mathbb{P}^{2d+1}(\mathbb{Z}/p^{k}\mathbb{Z})\setminus\bar{\mathcal{N}}_{p}^{(k)}=\{P\in\mathbb{P}^{2d+1}(\mathbb{Z}/p^{k}\mathbb{Z})\mathrel{\mathop{\ordinarycolon}}\ \mathfrak{r}(P)=0\ \text{in}\ \mathbb{Z}/p^{k}\mathbb{Z}\}.$
@@
The complement $(\bar{N}_{p}^{(k)})^{\prime}\mathrel{\mathop{\ordinarycolon}}=(\mathbb{Z}/p^{k}\mathbb{Z})^{2d+2}\setminus\bar{N}_{p}^{(k)}$ is the zero set of $\mathfrak{r}$ in $(\mathbb{Z}/p^{k}\mathbb{Z})^{2d+2}$ and the complement $(N_{p}^{(k)})^{\prime}\mathrel{\mathop{\ordinarycolon}}=\mathbb{Z}^{2d+2}\setminus N_{p}^{(k)}$ is the preimage of $(\bar{N}_{p}^{(k)})^{\prime}$ under the reduction map $\mathbb{Z}^{2d+2}\rightarrow(\mathbb{Z}/p^{k}\mathbb{Z})^{2d+2}$. As in Subsection 3.2, we warn the reader that neither $(\bar{N}_{p}^{(k)})^{\prime}$ is the affine lifting of $(\bar{\mathcal{N}}_{p}^{(k)})^{\prime}$ nor $(N_{p}^{(k)})^{\prime}$ is the affine lifting of $(\mathcal{N}_{p}^{(k)})^{\prime}$ (roughly speaking, this means taking the complement and taking the affine lifting do not commute to each other). In fact, the affine lifting of $(\bar{\mathcal{N}}_{p}^{(k)})^{\prime}$ is the set of points in $(\bar{N}_{p}^{(k)})^{\prime}$ having at least one entry that is a unit of $\mathbb{Z}/p^{k}\mathbb{Z}$. Similarly, the affine lifting of $(\mathcal{N}_{p}^{(k)})^{\prime}$ is the set of points in $(N_{p}^{(k)})^{\prime}$ having at least one entry that is not divisible by $p$. Therefore:
@@
In order to obtain an upper bound for $(\bar{N}_{p}^{(k)})^{\prime}$, we use the following result
@@
This result follows from [GKZ, Ch.3, Proposition 3.1 and Example 3.6]. ∎
@@
For the rest of this section, let $\mathfrak{s}$ denote one of the partial derivatives of $\mathfrak{r}$ (when necessary, which partial derivative will be specified precisely in later context). This is a homogeneous polynomial of degree 2d12d-1 and has degree d1d-1 in the variable for which we differentiate.
@@
$\{P\in\mathbb{F}_{p}^{2d+2}\mid\mathfrak{r}(P)=\mathfrak{s}(P)=0\}\leq\gamma_{1}p^{2d}$.
@@
Consider the algebra $\mathfrak{B}\mathrel{\mathop{\ordinarycolon}}=\mathbb{Q}[x_{0},\dots,x_{d},y_{0},\dots,y_{d}]/(\mathfrak{r},\mathfrak{s})$ and let $\bar{x}_{i}$ and $\bar{y}_{j}$ denote the image of $x_{i}$ and $y_{j}$ in $\mathfrak{B}$. Since $\mathfrak{r}$ is irreducible and does not divide $\mathfrak{s}$, the dimension of $\mathfrak{B}$ is $D\leq 2d$. By the Noether normalization lemma, there exist $X_{1},\ldots,X_{D}\in\mathfrak{B}$ that are algebraically independent over $\mathbb{Q}$ such that $\mathfrak{B}$ is integral over $\mathbb{Q}[X_{1},\dots,X_{D}]$. Thus there is a natural number $N$ such that the $\bar{x}_{i}$’s and $\bar{y}_{j}$’s are all integral over $\mathbb{Z}[1/N,X_{1},\dots,X_{D}]$. Denote by $\gamma_{1}^{\prime}$ the product of the degrees of the minimal polynomials of $\bar{x}_{i}$ and $\bar{y}_{j}$ over $\mathbb{Z}[1/N,X_{1},\dots,X_{D}]$.
Now suppose $p\nmid N$, and $(X_{1},\dots,X_{D})\in(\mathbb{Z}/p\mathbb{Z})^{D}$. Then each of the variables $x_{i}$ and $y_{j}$ satisfies a monic polynomial equation over $\mathbb{Z}/p\mathbb{Z}$. Let $g(z)\in\mathbb{Z}/p\mathbb{Z}[z]$ be their product and suppose that
$P=(A_{0},\dots,A_{d},B_{0},\dots,B_{d})\in(\mathbb{Z}/p\mathbb{Z})^{2d+2}$
is a point for which $\mathfrak{r}(P)=\mathfrak{s}(P)=0$. Then each of the coordinates $A_{i}$ and $B_{j}$ is a solution to $g(z)=0$. Since the degree of $g(z)$ is bounded by the constant $\gamma_{1}^{\prime}$ which is independent of the specialization $(X_{1},\dots,X_{D})$ and the prime $p$, we find that there are at most $(\gamma_{1}^{\prime})^{2d+2}$ choices of the point $P$ such that $\mathfrak{r}(P)=\mathfrak{s}(P)=0$. Thus for $p\nmid N$, the number of points $P$ is at most $(\gamma_{1}^{\prime})^{2d+2}p^{2d}$. Setting $\gamma_{1}\mathrel{\mathop{\ordinarycolon}}=\operatorname{max}\{(\gamma^{\prime})^{2d+2},N^{2}\}$, we find that the inequality in the statement of the proposition holds trivially for $p\mid N$. This finishes the proof. ∎
@@
Each element of $S_{1}$ lifts to $p^{2d+2}$ many elements of $(\mathbb{Z}/p^{2}\mathbb{Z})^{2d+2}$ and we have $S_{1}\leq\gamma_{1}p^{2d}$ by Proposition 4.13. Each element of $S_{2}$ lifts to $p^{2d+1}$ many elements $\beta\in(\mathbb{Z}/p^{2}\mathbb{Z})^{2d+2}$ such that $\mathfrak{r}(\beta)=0\bmod p^{2}$ by Lemma 4.15 and we have $S_{2}\leq 2dp^{2d+1}$ by Theorem 4.7. Overall, we have:
$(\mathcal{N}_{p}^{(2)})^{\prime}\leq(\gamma_{1}+2d)p^{4d+2}.$
@@
The set $W_{\Sigma}\mathrel{\mathop{\ordinarycolon}}=\left(\bigcap_{p\in\Sigma}G_{p}\right)\cap\left(\bigcap_{p\notin\Sigma}N_{p}^{(2d)}\right)$ is the affine lifting of $\mathcal{W}_{\Sigma}$. Recall the complement $(N_{p}^{(2d)})^{\prime}=\mathbb{Z}^{2d+2}\setminus N_{p}^{(2d)}$. We recall the notation at the beginning of Subsection 3.2. Let $Z$ be a positive real number such that $Z>\max\{p\in\Sigma\}$ and let
@@
$T_{>Z}(x)\mathrel{\mathop{\ordinarycolon}}=W_{>Z}^{\prime}(x)\setminus S_{>Z}(x).$
It follows from [Poonensquarefree, Lemma 5.1] that
@@
$h_{Q}(u)\mathrel{\mathop{\ordinarycolon}}=\mathfrak{r}(u,A_{1},\dots,A_{d},B_{0},\dots,B_{d-1},B_{d})\in\mathbb{Z}[u],$